hi
the areas of three faces of a rectangular box are 12 , 8 and 6. What is the volume of the box?
I have calculated surface area but am unsure how to calculate this problem

Answer by Alan3354(69443)

the areas of three faces of a rectangular box are 12 , 8 and 6. What is the volume of the box?
----------------
L*W = 12
L*H = 8
H*W = 6
---
L = 12/W
L*H = 8 --> 12H/W = 8 --> W = 3H/2
----
H*W = 6
H*(3H/2) = 6
3H^2 = 12
H = 2
----
W = 3H/2 = 3
----
L = 12/W = 4
----
Vol = 2*3*4 = 24 cubic units